# Chiral Heisenberg Gross-Neveu-Yukawa criticality: honeycomb vs. SLAC fermions

### v2

@@ III.2.3 Criticality
The question remains whether the effective long-range coupling alters the properties at the critical point which then would no longer be representative of the N=8{N=8} chiral Heisenberg universality class. First, we observe finite size single particle and spin gaps that vanish in the TDL (c.f. inset in Fig. 3, Fig. 8 and Fig. 9) consistent with gapless excitations in the two relevant sectors. Across criticality we further expect the fermionic Dirac spectrum to be replaced by its bosonic counterpart, such that the primary excitations share the same velocity [3, 71, 58]. In Fig. 9 we can see that at criticality and in contrast to Fig. 7 at strong coupling, the spin wave dispersion is isotropic, with an excitation velocity approaching the non-interacting Fermi velocity as indicated by the gray dashed line. This also provides an independent indicator that there is little to no renormalization of the Fermi velocity. Where the finite size dependence of the spin gaps at zero momentum is clear, for finite momenta this is less the case. Whether the shift to higher energies off the dashed line indicating the non-interacting Fermi velocity is a finite size effect or a persistent remnant of the dynamically induced long-range spin-exchange remains open. Again, we find no evidence for super-luminal magnon excitations. The recovered isotropy and similar velocities of the spin excitations compared to the fermionic single particle spectrum are in accordance with the expectations for the GNY-type transition.
@@
Figure 9: Close to the critical point the spin excitations appear isotropic and approximately reproduce the (non-interacting) Fermi velocity (dashed gray line, also cf. Fig. 5), albeit with an unclear finite size dependence for finite momenta.
@@
An additional worry would be, whether the dynamical long-range interactions preclude the fundamental necessary property of renormalizability. Evidence for the renormalizability is provided by dimensionless quantities, which remain finite with increasing system size, while the correlation length diverges as one approaches the critical point. The dimensionless ratios $R_{m^{2}}$ we introduce in the following Sec. IV fulfill this very property. Furthermore, we successfully employ a finite size scaling ansatz and are able to fit and collapse the finite size data accordingly. The working finite size scaling points towards the fact that the model is (at least) non-perturbatively renormalizable. We observe no pathological signatures up to the simulated lattice sizes. Note that in our investigation we do not need to take the continuum limit, but only the infinite volume limit, such that the lattice provides a natural, persistent ultraviolet cutoff.
@@
The dynamically induced long-range interactions does not appear to alter the critical properties. Indeed, long-range simulations of fermions on the honeycomb lattice subject to Coulomb interactions with a long-range part $V(r)\sim 1/r$} also have provided data that are approximately consistent with chiral Heisenberg Gross-Neveu universality [72, 13, 65] and is supported by long-range GNY models studied in [73].
@@ IV.1 Proximity to asymptotic scaling
It is arguable that many numerical investigations of critical phenomena, even when of the highest precision, fall short of the asymptotic scaling regime. This is the case for simulations that are not performed close enough to the critical point, or on insufficiently large lattices. Close, or large enough is determined by the crossover scale where the finite-size spectrum coincides with the chiral Heisenberg conformal field theory provided the linear extent of the simulation torus satisfies $L\gg|U-U_{c}|^{-\nu}$ [60]. In order to determine if a certain lattice model is closer than the another, one can directly compare complementary quantities, in order to identify favorable finite size scaling behavior.
@@
We first determine the critical point from the finite size intersections of pairs of CRs. We interpolate the data by second order polynomials (although predominantly linear) for each system size and then determine their intersections and the associated variance from a bootstrapping analysis. For the SLAC formulation we track the crossing points $R_{m^{2}}(u,L)=R_{m^{2}}(u,2L+1)$}, for the honeycomb we track $R_{m^{2}}(u,L)=R_{m^{2}}(u,2L)$}. We normalize their values with respect to the estimates of the critical points $U_{c}^{*}=U_{c}/5.98$} for the SLAC formulation and $U_{c}^{*}=U_{c}/3.77$} for the honeycomb lattice (c.f. Sec. IV.2), such that both should approach unity as $L^{-1/\nu-\min(\omega,2-z-\eta_{\phi})}$} for sufficiently large lattices [78]. While all dimensionless ratios should follow the same universal scaling behavior their approach can differ by nonuniversal factors. The comparison in Fig. 11 shows that the crossing points of the SLAC formulation significantly converge faster to the critical point despite the smaller lattice sizes. The figure further illustrates that all choices of the CRs approach the same critical point for large lattices.
Provided one has data close enough to the critical point, such that the scaling function can be approximated by a linear function, the slope of the CR $\partial R_{m^{2}}/\partial u\sim a_{1}L^{1/\nu}$} close to the critical point may be used as a measure of how fast the CR approaches the critical point and jumps from 0 to 1 in the TDL. The coefficient $a_{1}$ in Eq. (19) then determines how fast their slope diverges as $R_{m^{2}}$ becomes discontinuous as a function of the system size. Figure 12 illustrates the advantage of a larger $a_{1}$ in the SLAC formulation, i.e., the faster convergence with system size. The lines are linear fits to the finite size slopes on a double logarithmic scale and reflect the approximate system size dependence $\propto L^{1/\nu}$}  [^3] The difference of slopes at this scale is indistinguishable for different choices of $R_{m^{2}}^{{}_{(n_{1},n_{2})}}$ within the same lattice type.. This also provides a rough estimate for the critical exponent $\nu\approx 1$} which is shared among the two lattices.
Figure 10: Correlation ratios and their crossing points (black) as obtained from the intersections of polynomial fits to pairs $(L,2L+1)$ and $(L,2L)$ of the finite size data, for the SLAC formulation and the honeycomb, respectively.
@@
The two measures, the slope of the CRs close to criticality, which dominates the first term in Eq. (19), and the drift of the crossing points captured by scaling corrections in the second term show the SLAC formulation leads to more favorable finite size behavior.
IV.2 Finite size scaling analysis
We now apply the finite size scaling (FSS) ansatz for the finite size data of $R_{m^{2}}(u,L)$
@@ IV.2 Finite size scaling analysis
We consider the scaling ansatz without corrections to scaling first, i.e., $f_{1}^{R}=0$}. Including only system sizes $L\geq L_{\text{min}}$} in the analysis we obtain the distribution of the fit solutions in Fig. 13 with a single basin of solutions for fit parameters $\nu$ and $U_{c}$ (Fits from additional CRs are provided in App. B). The squared variances of the distributions for different choices of $L_{\text{min}}$ are indicated by the ellipses. Independently of the lattice type a correlation of the fit parameters is visible in the skewed distribution, such that smaller values of $\nu$ are more likely to be associated with larger values of $U_{c}$ and vice versa. We observe stable fits for the critical coupling and critical exponent $\nu$ across the different parameter sets. The fits do not significantly change for expansion orders $n_{\text{max}}\geq 1$, which supports that the considered coupling range is close enough to the critical point to observe dominant linear behavior of the CRs. If we allow for scaling corrections from irrelevant terms $f_{1}^{R}\neq 0$} we actually see that the additional degrees of freedom appear negligible for the SLAC fermions and tend to destabilize the fit as there is little drift in the finite size crossing points (c.f. Fig. 11 and App. B). The fits without scaling corrections allow for a direct comparison of convergence between the two lattice types. Where the SLAC fermions show little drift in the critical coupling $U_{c}$ and the exponent $\nu$, the estimates from the honeycomb lattice clearly move reflecting the crossing point analysis in Fig. 11. For large lattice sizes the estimates for $\nu$ coincide within error bars.
@@
The critical exponent of the correlations function associated with the order parameter $\eta_{\phi}$ can be extracted from the scaling the squared magnetization as a function of the CR $R_{m^{2}}$. This proves to be more stable than from the slope of $m^{2}$ as a function of system size, due to the larger amount of data to fit without the need of precise estimates of the critical point, thus eliminating the exponent $\nu$ from scaling. The FSS ansatz reduces to
@@
Figure 14: Fit of the finite size scaling ansatz to the squared magnetization in order to extract the bosonic anomalous dimensions $\eta_{\phi}$ for the SLAC formulation (top) and the honeycomb lattice (bottom). The insets illustrate the compatibility of the estimated exponents with the finite size decay behavior of the correlations at the largest distance.
As we track the evolution of the fits results with increasing minimum system size $L\geq L_{\text{min}}$} in Fig. 15, we see that $\eta_{\phi}$ approaches the same value for both lattices. Notably, the estimate from the SLAC formulation converges significantly faster, again suggestion improved finite size behavior of the SLAC simulations.
@@
We repeat the procedure to extract the anomalous dimension of the fermions $\eta_{\Psi}$ from the off-diagonal elements of the single-particle Green’s function at smallest lattice momentum $G_{ab}(\delta\mathbf{k})=\langle a^{\dagger}_{\delta\mathbf{k}}b^{\phantom{\dagger}}_{\delta\mathbf{k}}\rangle$} with the FSS ansatz
@@
The extracted the estimates for the critical exponents from the SLAC formulation read $\nu=1.02(3)$} for the correlation function, $\eta_{\phi}=0.73(1)$} for the anomalous dimension of the correlations associated with the order parameter and $\eta_{\psi}=0.09(1)$} for the anomalous dimension of the fermion correlations. The results coincide for the two lattice types once sufficiently large honeycomb lattices are considered. The consistency of the exponents between the two lattice types using the same scaling analysis favors the SLAC formulation for the superior finite size behavior. Note, the chiral Heisenberg exponents extracted here are incompatible with the universality of a long-range model O(3) spin models [80, 81].
@@ V Conclusion
The system at the critical point is at least non-perturbatively renormalizable and the extracted critical exponents presented here and now in several cases of simulations with a single Dirac cone in 2+1dd appear to be in accordance with existing estimates from local models, in particular in cases where we have exact results to compare with [33, 5, 27, 28, 24, 82]. Most importantly, they significantly reduce finite size effects, which are a major limiting factor in QMC simulations. Moreover, the SLAC formulation may be modified for a variety of low energy dispersions, which might be realized in local models, but where the physics in interacting systems is hidden at small energy scales which cannot be resolved in accessible lattice sizes of local Hamiltonians.
@@ Acknowledgements.
We thank F. F. Assaad for valuable discussions. This research was supported by the Austrian Science Fund FWF the SFB FoQuS (F-4018). We gratefully acknowledge the allocation of CPU time on the HPC infrastructure LEO of the University of Innsbruck and the Vienna Scientific Cluster VSC for allocation of CPU time.
@@ Appendix A Hopping amplitude of SLAC Fermions
For even $L$ system sizes we have the momenta $k=2\pi m/L$}, where $m=-M,\ldots,M-1\in\mathbb{N}$} with M=L/2{M=L/2} and we can use again use Eq. (23) to proceed with N=L/2+1{N=L/2+1} in the same way as for Eq. (25) to first compute
@@
Appendix B Additional correlations ratios & finite size scaling

### v3

@@ III.2.3 Criticality
The question remains whether the effective long-range coupling alters the properties at the critical point which then would no longer be representative of the N=8{N=8} chiral Heisenberg universality class. First, we observe finite-size single particle and spin gaps that vanish in the TDL (c.f. inset in Fig. 3, Fig. 8 and Fig. 9) consistent with gapless excitations in the two relevant sectors. Across criticality we further expect the fermionic Dirac spectrum to be replaced by its bosonic counterpart, such that the primary excitations share the same velocity [3, 70, 57]. In Fig. 9 we can see that at criticality and in contrast to Fig. 7 at strong coupling, the spin wave dispersion is isotropic, with an excitation velocity approaching the non-interacting Fermi velocity as indicated by the gray dashed line. This also provides an independent indicator that there is little to no renormalization of the Fermi velocity. Where the finite-size dependence of the spin gaps at zero momentum is clear, for finite momenta this is less the case. Whether the shift to higher energies off the dashed line indicating the non-interacting Fermi velocity is a finite-size effect or a persistent remnant of the dynamically induced long-range spin-exchange remains open. Again, we find no evidence for super-luminal magnon excitations. The recovered isotropy and similar velocities of the spin excitations compared to the fermionic single particle spectrum are in accordance with the expectations for the GNY-type transition.
@@
Figure 9: Close to the critical point the spin excitations appear isotropic and approximately reproduce the (non-interacting) Fermi velocity (dashed gray line, also cf. Fig. 5), albeit with an unclear finite-size dependence for finite momenta.
@@
An additional worry would be, whether the dynamical long-range interactions preclude the fundamental necessary property of renormalizability. Evidence for the renormalizability is provided by dimensionless quantities, which remain finite with increasing system size, while the correlation length diverges as one approaches the critical point. The dimensionless ratios $R_{m^{2}}$ we introduce in the following Sec. IV fulfill this very property. Furthermore, we successfully employ a finite-size scaling ansatz and are able to fit and collapse the finite-size data accordingly. The working finite-size scaling points towards the fact that the model is (at least) non-perturbatively renormalizable. We observe no pathological signatures up to the simulated lattice sizes. Note that in our investigation we do not need to take the continuum limit, but only the infinite volume limit, such that the lattice provides a natural, persistent ultraviolet cutoff.
@@
The dynamically induced long-range interactions does not appear to alter the critical properties. Indeed, long-range simulations of fermions on the honeycomb lattice subject to Coulomb interactions with a long-range part $V(r)\sim 1/r$} also have provided data that are approximately consistent with chiral Heisenberg Gross-Neveu universality [71, 13, 64] and are supported by long-range GNY models studied in [72].
@@ IV.1 Proximity to asymptotic scaling
It is arguable that many numerical investigations of critical phenomena, even when of the highest precision, fall short of the asymptotic scaling regime. This is the case for simulations that are not performed close enough to the critical point, or on insufficiently large lattices. Close, or large enough is determined by the crossover scale where the finite-size spectrum coincides with the chiral Heisenberg conformal field theory provided the linear extent of the simulation torus satisfies $L\gg|U-U_{c}|^{-\nu}$ [59]. In order to determine if a certain lattice model is closer than the another, one can directly compare complementary quantities, in order to identify favorable finite-size scaling behavior.
@@
We first determine the critical point from the finite-size intersections of pairs of CRs. We interpolate the data by second order polynomials (although predominantly linear) for each system size and then determine their intersections and the associated variance from a bootstrapping analysis. For the SLAC formulation we track the crossing points $R_{m^{2}}(u,L)=R_{m^{2}}(u,2L+1)$}, for the honeycomb we track $R_{m^{2}}(u,L)=R_{m^{2}}(u,2L)$}. We normalize their values with respect to the estimates of the critical points $U_{c}^{*}=U_{c}/5.98$} for the SLAC formulation and $U_{c}^{*}=U_{c}/3.77$} for the honeycomb lattice (c.f. Sec. IV.2), such that both should approach unity as $L^{-1/\nu-\min(\omega,2-z-\eta_{\phi})}$} for sufficiently large lattices [77]. While all dimensionless ratios should follow the same universal scaling behavior their approach can differ by nonuniversal factors. The comparison in Fig. 11 shows that the crossing points of the SLAC formulation significantly converge faster to the critical point despite the smaller lattice sizes. The figure further illustrates that all choices of the CRs approach the same critical point for large lattices.
Provided one has data close enough to the critical point, such that the scaling function can be approximated by a linear function, the slope of the CR $\partial R_{m^{2}}/\partial u\sim a_{1}L^{1/\nu}$} close to the critical point may be used as a measure of how fast the CR approaches the critical point and jumps from 0 to 1 in the TDL. The coefficient $a_{1}$ in Eq. (19) then determines how fast their slope diverges as $R_{m^{2}}$ becomes discontinuous as a function of the system size. Figure 12 illustrates the advantage of a larger $a_{1}$ in the SLAC formulation, i.e., the faster convergence with system size. The lines are linear fits to the finite-size slopes on a double logarithmic scale and reflect the approximate system size dependence $\propto L^{1/\nu}$}  [^3] The difference of slopes at this scale is indistinguishable for different choices of $R_{m^{2}}^{{}_{(n_{1},n_{2})}}$ within the same lattice type.. This also provides a rough estimate for the critical exponent $\nu\approx 1$} which is shared among the two lattices.
Figure 10: Correlation ratios and their crossing points (black) as obtained from the intersections of polynomial fits to pairs $(L,2L+1)$ and $(L,2L)$ of the finite-size data, for the SLAC formulation and the honeycomb, respectively.
@@
The two measures, the slope of the CRs close to criticality, which dominates the first term in Eq. (19), and the drift of the crossing points captured by scaling corrections in the second term show the SLAC formulation leads to more favorable finite-size behavior.
IV.2 finite-size scaling analysis
We now apply the finite-size scaling (FSS) ansatz for the finite-size data of $R_{m^{2}}(u,L)$
@@
We consider the scaling ansatz without corrections to scaling first, i.e., $f_{1}^{R}=0$}. Including only system sizes $L\geq L_{\text{min}}$} in the analysis we obtain the distribution of the fit solutions in Fig. 13 with a single basin of solutions for fit parameters $\nu$ and $U_{c}$ (Fits from additional CRs are provided in App. B). The squared variances of the distributions for different choices of $L_{\text{min}}$ are indicated by the ellipses. Independent of the lattice type a correlation of the fit parameters is visible in the skewed distribution, such that smaller values of $\nu$ are more likely to be associated with larger values of $U_{c}$ and vice versa. We observe stable fits for the critical coupling and critical exponent $\nu$ across the different parameter sets. The fits do not significantly change for expansion orders $n_{\text{max}}\geq 1$, which supports that the considered coupling range is close enough to the critical point to observe the dominant linear behavior of the CRs. If we allow for scaling corrections from irrelevant terms $f_{1}^{R}\neq 0$} we actually see that the additional degrees of freedom appear negligible for the SLAC fermions and tend to destabilize the fit as there is little drift in the finite-size crossing points (c.f. Fig. 11 and App. B). The fits without scaling corrections allow for a direct comparison of convergence between the two lattice types. Where the SLAC fermions show little drift in the critical coupling $U_{c}$ and the exponent $\nu$, the estimates from the honeycomb lattice clearly move reflecting the crossing point analysis in Fig. 11. For large lattice sizes the estimates for $\nu$ coincide within error bars.
@@
The critical exponent of the correlation function associated with the order parameter $\eta_{\phi}$ can be extracted from the scaling the squared magnetization as a function of the CR $R_{m^{2}}$. This proves to be more stable than from the slope of $m^{2}$ as a function of system size, due to the larger amount of data to fit without the need for precise estimates of the critical point, thus eliminating the exponent $\nu$ from scaling. The FSS ansatz reduces to
@@
Figure 14: Fit of the finite-size scaling ansatz to the squared magnetization in order to extract the bosonic anomalous dimensions $\eta_{\phi}$ for the SLAC formulation (top) and the honeycomb lattice (bottom). The insets illustrate the compatibility of the estimated exponents with the finite-size decay behavior of the correlations at the largest distance.
As we track the evolution of the fits results with increasing minimum system size $L\geq L_{\text{min}}$} in Fig. 15, we see that $\eta_{\phi}$ approaches the same value for both lattices. Notably, the estimate from the SLAC formulation converges significantly faster, again suggestion improved finite-size behavior of the SLAC simulations.
@@
We repeat the procedure to extract the anomalous dimension of the fermions $\eta_{\Psi}$ from the off-diagonal elements of the single-particle Green’s function at the smallest lattice momentum $G_{ab}(\delta\mathbf{k})=\langle a^{\dagger}_{\delta\mathbf{k}}b^{\phantom{\dagger}}_{\delta\mathbf{k}}\rangle$} with the FSS ansatz
@@
The extracted estimates for the critical exponents from the SLAC formulation read $\nu=1.02(3)$} for the correlation function, $\eta_{\phi}=0.73(1)$} for the anomalous dimension of the correlations associated with the order parameter and $\eta_{\psi}=0.09(1)$} for the anomalous dimension of the fermion correlations. The results coincide for the two lattice types once sufficiently large honeycomb lattices are considered. The consistency of the exponents between the two lattice types using the same scaling analysis favors the SLAC formulation for the superior finite-size behavior. Note, the chiral Heisenberg exponents extracted here are incompatible with the universality of a long-range model O(3) spin models [79, 80].
@@ V Conclusion
The system at the critical point is at least non-perturbatively renormalizable and the extracted critical exponents presented here and now in several cases of simulations with a single Dirac cone in 2+1dd appear to be in accordance with existing estimates from local models, in particular in cases where we have exact results to compare with [26, 5, 27, 28, 24, 81]. Most importantly, they significantly reduce finite-size effects, which are a major limiting factor in QMC simulations. Moreover, the SLAC formulation may be modified for a variety of low energy dispersions, which might be realized in local models, but where the physics in interacting systems is hidden at small energy scales that cannot be resolved in accessible lattice sizes of local Hamiltonians.
@@ Acknowledgements.
We thank F. F. Assaad for valuable discussions. We gratefully acknowledge the allocation of CPU time on the HPC infrastructure LEO of the University of Innsbruck and the Vienna Scientific Cluster VSC for the allocation of CPU time.
@@ Appendix A Hopping amplitude of SLAC Fermions
For even $L$ system sizes we have the momenta $k=2\pi m/L$}, where $m=-M,\ldots,M-1\in\mathbb{N}$} with M=L/2{M=L/2} and we can again use Eq. (23) to proceed with N=L/2+1{N=L/2+1} in the same way as for Eq. (25) to first compute
@@
Appendix B Additional correlations ratios & finite-size scaling